OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -21 and 567, respectively.

OpenStudy (campbell_st):

well a term in a geometric series can be found using \[a_{n} = a_{1} \times r^{n -1}\] so you need to find the 1st term and the common ratio using the information for the 5th term.... n = 2 you get \[-21 = a^{1} \times r^1\] now the 5th term \[567 = a_{1} \times r^{4}\] rewrite the 5th term to find r \[567 = a_{1} \times r \times r^3\] now substitute the 2nd term \[567 = -21 \times r^3\] now you can solve for the common ratio r... and then solve for a1, the 1st term hope it makes sense